How much work can an electric coffee dispenser do if it uses 25W for 

5min?


1
Expert's answer
2021-01-14T10:40:15-0500

By the definition of power, we have:


"P=\dfrac{W}{t},""W=Pt=25\ W\cdot5\ min\cdot \dfrac{60\ s}{1\ min}=7500\ J."

Answer:

"W=7500\ J."